# Sample-Efficient Reinforcement Learning from Human Feedback via Information-Directed Sampling

### v1

@@ V-2 Regret bounds for Approximate-IDS
$\leq\frac{\lambda\epsilon(1-2\log\beta)}{2}+\mathbb{E}_{\pi_{\mathrm{app}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})-r_{h}(s_{h},a_{h})\bigg{]}$
$\leq\frac{\lambda\epsilon(1-2\log\beta)}{2}+\frac{\lambda}{2}\cdot\mathbb{I}_{t}^{\pi_{\mathrm{app}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right),$ (23)
where the first inequality is due to Eqn. (V-2), the second inequality is due to Proposition 1, and the last inequality is due to Lemma 5. Taking expectation in Eqn. (V-2) with respect to $\mathcal{D}_{t}$ and then summing over $t\in[T]$, we obtain
$BR_{T}(\pi_{\mathrm{app}})-BR_{T}(\pi_{\mathrm{IDS}})\leq\frac{\lambda\epsilon(1-2\log\beta)T}{2}+\frac{\lambda}{2}\log(K(\epsilon)),$ (24)
where we use the same trick in Eqn. (A-A) to derive $\log(K(\epsilon))$ as an upper bound for $\sum_{t=1}^{T}\mathbb{I}_{t}^{\pi_{\mathrm{app}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$. Finally, plugging the upper bound for $BR_{T}(\pi_{\mathrm{IDS}})$ (Eqn. (A-A)) into Eqn. (24) and taking $\lambda=\sqrt{\alpha^{2}TH/2\log(K(\epsilon))}$ yields the proof of Theorem 4. ∎
VI Conclusion
@@ A-A Proof of Theorem 1
$BR_{T}(\pi_{\mathrm{IDS}})\leq\alpha\sqrt{TH\log(K(\epsilon))}+T\epsilon+T_{0},$ (25)
@@
$BR_{T}(\pi_{\text{IDS}})=\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{\mathcal{E}\sim\mathbb{P}(\cdot|\mathcal{D}_{t})}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi^{t}}^{\mathcal{E}}(s_{1}^{t})\right]\right],$ (26)
@@
Step 1. Reduce $BR_{T}(\pi_{\mathrm{IDS}})$ to the surrogate environment, and convert $BR_{T}(\pi_{\mathrm{IDS}})$ into $BR_{T}(\pi_{\mathrm{TS}})$. By Lemma 2 and the optimality of $\pi_{\mathrm{IDS}}$, we have
$BR_{T}(\pi_{\mathrm{IDS}})=\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{\mathcal{E}\sim\mathbb{P}(\cdot|\mathcal{D}_{t})}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathrm{IDS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]\right]$
$=\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathrm{IDS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]-\epsilon-\frac{\lambda}{2}\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]+\frac{\lambda}{2}\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]+T\epsilon$
$\leq\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]-\frac{\lambda}{2}\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]+\frac{\lambda}{2}\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]+T\epsilon.$ (27)
For the first term in Eqn. (A-A), using the basic fact that $A-\lambda B/2\leq A^{2}/2\lambda B$ for $B,\lambda\geq 0$, we have
$\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]-\frac{\lambda}{2}\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\leq\frac{1}{2\lambda}\frac{\left(\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]\right)^{2}}{\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)}\triangleq\frac{1}{2\lambda}\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}.$ (28)
@@
Let $\zeta$ be a discrete random variable taking values in $\{1,...,K(\epsilon)\}$ such that $\zeta=k$ if and only if $\mathcal{E}\in\Theta_{k}^{\epsilon}$. From the construction of the surrogate environment (Eqn. (10)), the distribution of ~$\tilde{\mathcal{E}}_{t}^{*}$ depend on $\mathcal{E}$ only through $\zeta$, i.e., ~$\tilde{\mathcal{E}}_{t}^{*}$ and $\mathcal{E}$ are independent conditioning on $\zeta$.
For the second term in Eqn. (A-A), we have
$\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]$ $\leq\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\zeta;(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]$
$=\mathbb{I}(\zeta;\mathcal{D}_{T+1})$
$\leq\mathbb{H}(\zeta)$
$\leq\log(K(\epsilon)).$ (29)
@@
$BR_{T}(\pi_{\mathrm{IDS}})\leq\frac{1}{2\lambda}\mathbb{E}\left[\sum_{t=1}^{T}\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}\right]+\frac{\lambda}{2}\log(K(\epsilon))+T\epsilon.$ (30)
@@
For convenience of analysis, we assume that $d_{h,\pi}^{\mathcal{E}}(s,a)$ is measurable and upper bounded for all $\pi,\mathcal{E},s,a,h$. Recall that, the mean environment ¯$\bar{\mathcal{E}}_{t}$ is defined to satisfy $P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[P_{h}^{\mathcal{E}}(\cdot|s,a)]$ and $R_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[R_{h}^{\mathcal{E}}(\cdot|s,a)]$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$. By the definition of $d_{h,\pi}^{\mathcal{E}}$, the following equality also holds: $d_{h,\pi}^{\bar{\mathcal{E}}_{t}}(s,a)=\mathbb{E}_{t}[d_{h,\pi}^{\mathcal{E}}(s,a)]$. One important property of the mean environment is that the posterior mean of the surrogate environment $\mathbb{E}_{t}[\tilde{\mathcal{E}}_{t}^{*}]$ coincides with that of the whole environment ¯$\bar{\mathcal{E}}_{t}$. To check this, using the property of conditional expectation:
$\mathbb{E}_{t}[\tilde{\mathcal{E}}_{t}^{*}]$ $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}[\tilde{\mathcal{E}}_{t}^{*}|\mathcal{E}\in\Theta_{k}^{\epsilon}]$
$=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}[\tilde{\mathcal{E}}_{k,t}^{*}]$
$=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\tilde{\mathcal{E}}_{k,t}^{*}$
$=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}[\mathcal{E}|\mathcal{E}\in\Theta_{k}^{\epsilon}]=\bar{\mathcal{E}}_{t}.$ (31)
@@
$\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)\triangleq\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\tilde{\mathcal{E}}_{t}^{*}})(\cdot|s,a)}\left[r^{\prime}+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s^{\prime})\right]-\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s,a)}\left[r^{\prime}+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s^{\prime})\right].$ (32)
Now we are ready to give an upper bound for $\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}$. We hope to use Lemma 4 to rewrite the numerator
$\left(\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]\right)^{2}.$
However, Lemma 4 can only be applied to handle the difference between two value functions with the same policy and different environments, while in $V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})$ and $V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})$, the environments are the same and the policies are different. For the purpose of “unifying” the policy, we use Eqn. (31) and note that $\pi_{\mathrm{TS}}$ is independent of ~$\tilde{\mathcal{E}}_{t}^{*}$, yielding
$\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]=\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right].$
Furthermore, conditioned on $\mathcal{D}_{t}$, $\pi_{\text{TS}}^{t}$ and $\pi_{\mathcal{E}}^{*}$ are identically distributed, and are both independent of ¯$\bar{\mathcal{E}}_{t}$. This implies

### v3

@@ 5.2 Regret bounds for Approximate-IDS
$\leq\frac{\lambda\epsilon(1+\log(B/\beta))}{2}+\mathbb{E}_{\pi_{\mathrm{app}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})-r_{h}(s_{h},a_{h})\bigg{]}$
$\leq\frac{\lambda\epsilon(1+\log(B/\beta))}{2}+\frac{\lambda}{2}\cdot\mathbb{I}_{t}^{\pi_{\mathrm{app}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right),$ (5.7)
where the first inequality is due to Eqn. (5.2), the second inequality is due to Proposition 5.1, and the last inequality is due to Lemma C.2. Taking expectation in Eqn. (5.2) with respect to $\mathcal{D}_{t}$ and then summing over $t\in[T]$, we obtain
$BR_{T}(\pi_{\mathrm{app}})-BR_{T}(\pi_{\mathrm{IDS}})\leq\frac{\lambda\epsilon(1+\log(B/\beta))T}{2}+\frac{\lambda}{2}\log(K(\epsilon)),$ (5.8)
where we use the same trick in Eqn. (A.5) to derive $\log(K(\epsilon))$ as an upper bound for $\sum_{t=1}^{T}\mathbb{I}_{t}^{\pi_{\mathrm{app}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)$. Finally, plugging the upper bound for $BR_{T}(\pi_{\mathrm{IDS}})$ (Eqn. (A.1)) into Eqn. (5.8) and taking $\lambda=\sqrt{\alpha^{2}TH/2\log(K(\epsilon))}$ yields the proof of Theorem 5.2. ∎
5.3 Regret Analysis With $\ell_{1}$-distance
Using the standard $\ell_{1}$-distance for both environment partitioning and surrogate environment construction, rather than the proposed $\ell_{g}$-distance, would necessitate significantly stronger and less realistic assumptions, and result in worse theoretical guarantees. To ensure the validity of Proposition 5.1, the following stronger assumptions should be additionally imposed.
Assumption 5.3.
For any $P\in\Theta$, there exist $\beta,B>0$ such that
$\beta\leq\inf_{o,x}\{P(x|o)\}\leq\sup_{o,x}\{P(x|o)\}\leq B.$ (5.9)
This assumption excludes zero probability densities for transitions/rewards, limiting its practical applicability significantly. Under this assumption, Proposition 5.1 can be reformulated as follows.
Proposition 5.4.
For any $\epsilon$-value partition $\{\Theta_{k}^{\epsilon}\}_{k=1}^{K(\epsilon)}$ (partitioned by $\ell_{1}$-distance) and the surrogate environment $\widetilde{\mathcal{E}}_{t}^{*}$ (constructed by $\ell_{1}$-distance ), we have
$\bigg{|}\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}r^{\prime}_{h}(s_{h},a_{h})\bigg{]}-\mathbb{E}_{\pi}^{\bar{\mathcal{E}}_{t}}\bigg{[}\sum_{h=1}^{H}\bar{r}_{h}(s_{h},a_{h})\bigg{]}\bigg{|}\leq\frac{\lambda}{2}\epsilon(1+\frac{B}{\beta}).$ (5.10)
The proof of Proposition 5.4 can be found in Appendix A.5. Using Proposition 5.4, the Bayesian regret bound for Algorithm 2 under $\ell_{1}$-distance is of order
$\widetilde{O}(\frac{1}{\beta}\sqrt{H^{3}T}).$
In comparison, the regret bound derived using $\ell_{g}$ is of order
$\widetilde{O}((\log\frac{1}{\beta})\sqrt{H^{3}T}).$
Although we have also derived a regret bound of sublinear dependence on $H$ and $T$ using the $\ell_{1}$-distance, this bound includes a factor of $\frac{1}{\beta}$. When $\beta$ is small, the bound becomes excessively large. Moreover, the underlying assumption of this bound contradicts real-world scenarios, as common MDPs permit zero transition and reward probabilities in practical settings—such as the boundaries in Go or obstacles in robotic navigation. In light of these limitations, our proposed $\ell_{g}$-based approach demonstrates significant superiority.
6 Conclusion
@@ Appendix A Proofs of Theorem and Proposition
$BR_{T}(\pi_{\mathrm{IDS}})\leq\alpha\sqrt{TH\log(K(\epsilon))}+T\epsilon+T_{0},$ (A.1)
@@
$BR_{T}(\pi_{\text{IDS}})=\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{\mathcal{E}\sim\mathbb{P}(\cdot|\mathcal{D}_{t})}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi^{t}}^{\mathcal{E}}(s_{1}^{t})\right]\right],$ (A.2)
@@
Step 1. Reduce $BR_{T}(\pi_{\mathrm{IDS}})$ to the surrogate environment, and convert $BR_{T}(\pi_{\mathrm{IDS}})$ into $BR_{T}(\pi_{\mathrm{TS}})$. By Lemma 4.9 and the optimality of $\pi_{\mathrm{IDS}}$, we have
$BR_{T}(\pi_{\mathrm{IDS}})$ $=\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{\mathcal{E}\sim\mathbb{P}(\cdot|\mathcal{D}_{t})}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathrm{IDS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]\right]$ (A.3)
$=\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathrm{IDS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]-\epsilon-\frac{\lambda}{2}\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]$
$\qquad+\frac{\lambda}{2}\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]+T\epsilon$
$\overset{(a)}{\leq}\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\mathcal{E}}(s_{1}^{t})\right]-\epsilon-\frac{\lambda}{2}\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]$
$\qquad+\frac{\lambda}{2}\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]+T\epsilon$
$\overset{(b)}{\leq}\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]-\frac{\lambda}{2}\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]$
$\qquad+\frac{\lambda}{2}\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]+T\epsilon,$
where (a) uses the optimality of $\pi_{\mathrm{IDS}}^{t}$, (b) uses Lemma 4.9.
For the first term in Eqn. (A.3), using the basic fact that $A-\lambda B/2\leq A^{2}/2\lambda B$ for $B,\lambda\geq 0$, we have
$\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]-\frac{\lambda}{2}\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\leq\frac{1}{2\lambda}\frac{\left(\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]\right)^{2}}{\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)}\triangleq\frac{1}{2\lambda}\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}.$ (A.4)
@@
Let $\zeta$ be a discrete random variable taking values in $\{1,...,K(\epsilon)\}$ such that $\zeta=k$ if and only if $\mathcal{E}\in\Theta_{k}^{\epsilon}$. From the construction of the surrogate environment (Eqn. (4.9)), the distribution of $\widetilde{\mathcal{E}}_{t}^{*}$ depend on $\mathcal{E}$ only through $\zeta$, i.e., $\widetilde{\mathcal{E}}_{t}^{*}$ and $\mathcal{E}$ are independent conditioning on $\zeta$.
For the second term in Eqn. (A.3), we have
$\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]\leq\sum_{t=1}^{T}\mathbb{E}_{\mathcal{D}_{t}}\left[\mathbb{I}_{t}^{\pi_{\mathrm{IDS}}^{t}}\left(\zeta;(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)\right]=\mathbb{I}(\zeta;\mathcal{D}_{T+1})\leq\mathbb{H}(\zeta)\leq\log(K(\epsilon)),$ (A.5)
@@
$BR_{T}(\pi_{\mathrm{IDS}})\leq\frac{1}{2\lambda}\mathbb{E}\left[\sum_{t=1}^{T}\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}\right]+\frac{\lambda}{2}\log(K(\epsilon))+T\epsilon.$ (A.6)
@@
For convenience of analysis, we assume that $d_{h,\pi}^{\mathcal{E}}(s,a)$ is measurable and upper bounded for all $\pi,\mathcal{E},s,a,h$. Recall that, the mean environment $\bar{\mathcal{E}}_{t}$ is defined to satisfy $P_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[P_{h}^{\mathcal{E}}(\cdot|s,a)]$ and $R_{h}^{\bar{\mathcal{E}}_{t}}(\cdot|s,a)=\mathbb{E}_{t}[R_{h}^{\mathcal{E}}(\cdot|s,a)]$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$. By the definition of $d_{h,\pi}^{\mathcal{E}}$, the following equality also holds: $d_{h,\pi}^{\bar{\mathcal{E}}_{t}}(s,a)=\mathbb{E}_{t}[d_{h,\pi}^{\mathcal{E}}(s,a)]$. One important property of the mean environment is that the posterior mean of the surrogate environment $\mathbb{E}_{t}[\widetilde{\mathcal{E}}_{t}^{*}]$ coincides with that of the whole environment $\bar{\mathcal{E}}_{t}$. To check this, using the property of conditional expectation:
$\mathbb{E}_{t}[\widetilde{\mathcal{E}}_{t}^{*}]$ $=\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}[\widetilde{\mathcal{E}}_{t}^{*}|\mathcal{E}\in\Theta_{k}^{\epsilon}]$
$\overset{(a)}{=}\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}[\widetilde{\mathcal{E}}_{k,t}^{*}]$
$\overset{(b)}{=}\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\widetilde{\mathcal{E}}_{k,t}^{*}$
$\overset{(c)}{=}\sum_{k=1}^{K}\mathbb{P}(\mathcal{E}\in\Theta_{k}^{\epsilon})\cdot\mathbb{E}_{t}[\mathcal{E}|\mathcal{E}\in\Theta_{k}^{\epsilon}]=\bar{\mathcal{E}}_{t},$ (A.7)
where (a) comes from the definition of $\widetilde{\mathcal{E}}_{t}^{*}$ and $\widetilde{\mathcal{E}}_{k,t}^{*}$, (b) and (c) uses the following fact
$\mathbb{E}_{t}[\widetilde{\mathcal{E}}_{k,t}^{*}]=\mathbb{E}_{t}[\mathbb{E}_{t}[\mathcal{E}|\mathcal{E}\in\Theta_{k}^{\epsilon}]]=\mathbb{E}_{t}[\mathcal{E}|\mathcal{E}\in\Theta_{k}^{\epsilon}]=\widetilde{\mathcal{E}}_{t}^{*}$
@@
$\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)\triangleq\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}})(\cdot|s,a)}\left[r^{\prime}+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})\right]-\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s,a)}\left[r^{\prime}+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})\right].$ (A.8)
Now we are ready to give an upper bound for $\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}$. We hope to use Lemma C.1 to rewrite the numerator
$\left(\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]\right)^{2}.$
However, Lemma C.1 can only be applied to handle the difference between two value functions with the same policy and different environments, while in $V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})$ and $V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})$, the environments are the same and the policies are different. For the purpose of “unifying” the policy, we use Eqn. (A.7) and note that $\pi_{\mathrm{TS}}$ is independent of $\widetilde{\mathcal{E}}_{t}^{*}$, yielding
$\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]=\mathbb{E}_{t}\left[V_{1,\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right].$
Furthermore, conditioned on $\mathcal{D}_{t}$, $\pi_{\text{TS}}^{t}$ and $\pi_{\mathcal{E}}^{*}$ are identically distributed, and are both independent of $\bar{\mathcal{E}}_{t}$. This implies